Find the third Taylor polynomial of the function f(x,y)= =1+5xy+3²y at (1,2)


1
Expert's answer
2021-05-19T18:15:06-0400

Given "f(x,y)=1-5xy+3^2y=1-5xy+9y"


"f(1,2)=1-5(1)(2)+9(2)=1-10+18=9"


"f_x=-5y\\

f_x(1,2)=-10


\\
f_y=-5x+9 

 
\\
f_y(1,2)=-5(1)+9=4




\\
f_{xx}=0\\

f_{yy}=0
\\"



"f(x,y)=f(a,b)+(x-1)f_x(1,2)+(y-2)f_y+\dfrac{f_{xx}\times (x-1)^2}{2!}+...."


"=9+(x-1)\times -10+(y-2)\times 4+0+...\\[9pt]=9-10x+10+4y-8\\[9pt]=4y-10x+11"